Week of 190211

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Alberto (Monitoring), Borja (Chair, Monitoring), Gavin (Computing), Julia (WLCG), Maarten (ALICE), Marian (Network), Pter (ATLAS)
  • remote: Andrew (NLT1), Christoph (CMS), Di (TRIUMF), Dave (FNAL), John (RAL), Jose (PIC), Marcelo (CNAF), Ville (NDGF), Zoltan (LHCb)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Production stable with low number of errors and additional CERN P1 resources
    • Influxdb down on Sunday for 16 hours - ATLAS DDM Transfer monitoring source
    • DPM - problems after upgrade to DOME

Borja acknowledged the issue with InfluxDB producing a downtime for the new ATLAS DDM Transfers monitoring, issue required a manual restart and was done in the same day. It has been followed up with the DBOD team at CERN.

There was some discussion about the upgrade of DPM with some sites not being updated correctly, CMS agreed. Maarten explained the recipes for the admins are still to be improved but in the meanwhile tickets can be opened exposing the issues.

  • CMS reports ( raw view) -
    • Very good CPU utilization ~260k cores (passed over 300k at some occasions)
      • Usual mix of ~75% production, 25% analysis
    • No major issues
    • Issue with SAM test submission since the weekend GGUS:139606

Marian commented on the SAM test submission issue, seems to be an issue with HTCondor itself, not CERN related. Similar issue happened already some weeks ago and was fixed by itself. Maarten asked if it has been checked with Condor developers, it's not been yet. It will be followed up with them, client version was mentioned as one of the possible causes.

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity
      • User jobs and MC product
      • Stripping s35
    • Sites Issues

Sites / Services round table:

  • ASGC: NC
  • BNL: NC
  • CNAF: Data transfer problem from LHCb (GGUS:139608). Problem seems solved.
  • EGI: NC
  • FNAL:
    • Issue transferring data to the EOS instance, seems it has been taken care of already
    • Lost one cache controller in T1 due to hardware failure, around 3K files were lost none of them containing raw data, so no data loss.
  • IN2P3: NTR
  • JINR: NTR
  • KISTI: NC
  • KIT: NTR
  • NDGF: dCache master servers will be upgraded to version 5 tomorrow. Hopefully only a short break in services: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=26727
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: Set up new end point for AliceTape. Otherwise there is nothing to report.
  • TRIUMF: NTR

  • CERN computing services:
    • OTG:0047300: HTCondor is now at ~30% CC7, proceeding with draining to get to 50% as per:
      • end March 2019: 50% public/grid will be CC7
      • 2nd April 2019: lxplus.cern.ch alias change to CC7 (lxplus6 service will remain accessible on lxplus6.cern.ch), Default HTCondor target change to CC7 for local submission.
      • early June 2019: remainder of capacity will have been migrated.
      • (Dedicated shares handled separately.)
    • OTG:0048002 ce511, ce512, ce513, ce514 now default to CC7.
  • CERN storage services:
    • OTG:0048008 Migration of EOSLHCB to new backend: 2 hours downtime (see also GOCDb )
  • CERN databases: NC
  • GGUS: NTR
  • Monitoring:
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2019-02-11 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback